1. Greg bought a laptop for P28000 and spent P2500 on its spares. He later sold it at P25000. How much is greg's loss?


2. Alden went to bicol and bought 25 packs of pili nuts for P875 with 10% discount. When he got back to manila, he sold 15 packs for P675, and the rest at P40 each. How much profit did alden gain?


3. Arman sold a watch to his friend at a loss 5%. If his friend paid P650 for it, find the original cost of the watch.


4. A mobile phone was for P4800. Find the cost of the mobile phone if the profit is 20% .


5. A shopkeeper buys digital camera at P5400. Due to the fall in the market, he was forced to sell them at loss of 15%. Find the selling price for each digital camera.

1
Expert's answer
2021-11-08T16:35:30-0500

1) Loss= Selling price - Buying price - Related costs

Loss = 25000-28000-2500

Loss= P5500

2) Buying price after discount = 875*90%

Buying price after discount = P787.50

Selling price = 15*675 = P10125

Other packets = 10*40 = P400

Total Selling price = 10125+400 = P10525

Profit = Selling price - Buying Price

Profit= 10525-787.50

Profit = 9737.50

3) Original Price = 100*650/95

Original Price = P684.20

4) Cost of mobile phone = 100*4800/120

Cost of the Mobile =P4000

5) Selling price = 85/100*5400

Selling price at a loss = P4590